# The Angular Momentum Penrose Inequality

### v3

@@ 6.2 The AM-AMO Functional
  • Area: $A(t):=\int_{\Sigma_{t}}dA_{\tilde{g}}$
[…]
  • Willmore functional: $W(t):=\int_{\Sigma_{t}}H^{2}\,dA_{\tilde{g}}$ (the unnormalized Willmore energy)
  • Hawking mass: $m_{H}(t):=\sqrt{\frac{A(t)}{16\pi}}\left(1-\frac{W(t)}{16\pi}\right)$, defined when $W(t)\leq 16\pi$
@@
$m_{H,J}(t):=\sqrt{m_{H}^{2}(t)+\frac{4\pi J^{2}}{A(t)}}=\sqrt{\frac{A(t)}{16\pi}\left(1-\frac{W(t)}{16\pi}\right)^{2}+\frac{4\pi J^{2}}{A(t)}},$ (46)
@@
Well-definedness: For sub-extremal surfaces with $A(t)\geq 8\pi|J|$ (ensured by Theorem 7.2), the argument of the outer square root is non-negative. The Willmore bound $W(t)\leq 16\pi$ is established in Lemma 6.10 below.
@@
  • The first correction term satisfies $2\int_{\Sigma}R_{\tilde{g}}\,dA\leq 0$;
  • The Ricci term requires more care. Using the contracted Gauss equation:
    $R_{\tilde{g}}=R_{\Sigma}+|A|^{2}-H^{2}+2\mathrm{Ric}_{\tilde{g}}(\nu,\nu).$
    Rearranging: $\mathrm{Ric}_{\tilde{g}}(\nu,\nu)=\frac{1}{2}(R_{\tilde{g}}-R_{\Sigma}-|A|^{2}+H^{2})$.
@@
$\frac{1}{16\pi}\int_{\Sigma_{t}}H^{2}\,dA\leq 1\quad\Longleftrightarrow\quad W(t)=\int_{\Sigma_{t}}H^{2}\,dA\leq 16\pi.$
Step 8: Verification at boundary and infinity.
  • At $t=0$ (MOTS): The MOTS $\Sigma$ has $H_{\tilde{g}}|_{\Sigma}=0$ (minimal in the conformal metric, see Lemma 8.4), so $W(0)=0<16\pi$.
  • At $t\to 1$ (infinity): Large coordinate spheres of radius $R$ have $H\approx 2/R$, area $\approx 4\pi R^{2}$, so $W\approx(4/R^{2})(4\pi R^{2})=16\pi(1-O(1/R))$.
  • For intermediate $t$: The monotonicity $m_{H}^{\prime}(t)\geq 0$ implies $(1-W(t)/(16\pi))$ remains non-negative throughout.
The bound $W(t)\leq 16\pi$ is therefore a consequence of the AMO monotonicity framework, not an independent assumption. ∎
@@
[…]
  • $W(t)/(16\pi)=0$ when $H\equiv 0$ (minimal surfaces);
  • $W(t)/(16\pi)\in[0,1]$ for topological 2-spheres with non-negative scalar curvature ambient manifolds;
  • $W(t)/(16\pi)\to 1^{-}$ as $t\to 1$ (large coordinate spheres).
Some references define the “Willmore deficit” as $\mathcal{W}:=W/(16\pi)$. In this paper, when we write “(1W)(1-W)” in formulas, we always mean $(1-W/(16\pi))$, not (1W)(1-W) literally. This convention is consistent with the Hawking mass formula $m_{H}=\sqrt{A/(16\pi)}(1-W/(16\pi))$.
@@
$\frac{W(t)}{16\pi}=\frac{1}{16\pi}\int_{\Sigma_{t}}H^{2}\,d\sigma\approx\frac{1}{16\pi}\cdot 4\pi R^{2}\cdot\frac{4}{R^{2}}=1-O(R^{-1}).$
@@ 6.3 Angular Momentum Conservation
Answer: The key insight is the separation of roles:
  • The underlying smooth manifold is the same: $M=\bar{M}=\tilde{M}$ as topological spaces (the Jang and conformal constructions are diffeomorphisms, not changes of the underlying manifold).
  • The level sets $\Sigma_{t}=\{u=t\}$ are embedded submanifolds of $M$, defined using $\tilde{g}$ but living in the same $M$ where $(g,K)$ are defined.
  • The Komar 1-form $\alpha_{J}=\frac{1}{8\pi}K(\eta,\cdot)^{\flat}_{g}$ is a well-defined 1-form on $M$, independent of any choice of Riemannian metric.
  • The integral $J(\Sigma_{t})=\int_{\Sigma_{t}}\star_{g}\alpha_{J}$ is computed using the Hodge dual with respect to the physical metric $g$, not the conformal metric $\tilde{g}$.
Thus $J(\Sigma_{t})$ is a well-defined quantity: integrate the fixed 2-form $\star_{g}\alpha_{J}$ (determined by $(g,K)$) over the surface $\Sigma_{t}$ (located using $\tilde{g}$).
@@
Answer: We do not claim that $d^{\dagger}_{\tilde{g}}\alpha_{J}=0$. Instead:
  • Co-closedness is established for the physical metric: $d^{\dagger}_{g}\alpha_{J}=0$.
  • This is equivalent to: $d(\star_{g}\alpha_{J})=0$, i.e., $\star_{g}\alpha_{J}$ is a closed 2-form.
  • The exterior derivative $d$ is metric-independent—it is a purely topological operation.
  • Therefore $d(\star_{g}\alpha_{J})=0$ holds on the smooth manifold $M$ regardless of which metric is used to parametrize surfaces.
The conservation law is a consequence of Stokes’ theorem applied to the closed 2-form $\star_{g}\alpha_{J}$, not a conformal invariance statement.
@@
Answer: Yes. The constructions preserve axisymmetry:
  • The Jang equation with axisymmetric boundary conditions yields an axisymmetric solution $f$, so $\mathcal{L}_{\eta}\bar{g}=0$ where $\bar{g}=g+df\otimes df$.
  • The AM-Lichnerowicz equation with axisymmetric data yields an axisymmetric conformal factor $\phi$, so $\mathcal{L}_{\eta}\tilde{g}=0$ where $\tilde{g}=\phi^{4}\bar{g}$.
  • Therefore $\eta$ remains a Killing field for $\tilde{g}$, and the $p$-harmonic flow respects the symmetry.
However, this is not needed for conservation: even if $\eta$ were not Killing for $\tilde{g}$, the closed form $\star_{g}\alpha_{J}$ would still have constant flux through homologous surfaces.
@@
Answer: The Jang manifold $\bar{M}$ has a cylindrical end $\mathcal{C}\cong[0,\infty)\times\Sigma$. Key points:
  • The Komar 1-form $\alpha_{J}$ extends smoothly to the cylindrical end (it is defined from $(g,K)$, which are smooth).
  • The 2-form $\star_{g}\alpha_{J}$ is closed throughout $M$, including the cylindrical region.
  • The level sets $\Sigma_{t}$ for $t$ near 0 may approach the MOTS $\Sigma=\Sigma_{0}$, but remain in a region where $\star_{g}\alpha_{J}$ is defined.
  • The flux $\int_{\Sigma_{t}}\star_{g}\alpha_{J}$ is continuous in $t$, even as $t\to 0$, by dominated convergence.
The boundary term at the cylindrical end vanishes by the asymptotic analysis in Lemma 5.9.
@@
Theorem 6.14 (Angular Momentum Conservation—Topological).
@@
Step (ii): Coordinate calculation. Let $(y^{1},y^{2})$ be local coordinates on $\Sigma$ and extend to coordinates $(y^{1},y^{2},n)$ on $U$ where $n$ is a coordinate transverse to $\Sigma$ with $\Sigma=\{n=0\}$. In these coordinates:
  • The $\gamma$-unit normal is $\nu_{\gamma}=\frac{1}{|\partial_{n}|_{\gamma}}\partial_{n}+(\text{tangential corrections})$.
  • The area element is $d\sigma_{\gamma}=|\partial_{n}|_{\gamma}\sqrt{\det\gamma_{AB}}\,dy^{1}\wedge dy^{2}$ where $\gamma_{AB}$ is the induced metric on $\Sigma$.
  • The contraction $\alpha_{J}(\nu_{\gamma})=\frac{1}{|\partial_{n}|_{\gamma}}(\alpha_{J})_{n}+(\text{tangential terms})$.
The product gives:
@@
The key observation is that the Komar 1-form $\alpha_{J}=\frac{1}{8\pi}K(\eta,\cdot)^{\flat}$ is defined on the physical manifold, but we integrate it over surfaces $\Sigma_{t}$ that are level sets in the conformal picture. This is valid because:
  1. 1.
    The underlying smooth manifold $M$ is the same; only the metric changes.
  2. 2.
    The level sets $\Sigma_{t}\subset M$ are well-defined submanifolds independent of which metric we use.
  3. 3.
    The 1-form $\alpha_{J}$ and its exterior derivative $d\alpha_{J}$ are tensorial operations that commute with pullback to any submanifold.
  4. 4.
    The integral $\int_{\Sigma_{t}}\star_{g}\alpha_{J}$ is computed using the physical metric $g$ for the Hodge dual, making it independent of $\tilde{g}$.
@@
Metric-independence of the Komar integral. We now make explicit which quantities use which metric. Define:
  • $\nu_{\tilde{g}}:=\nabla_{\tilde{g}}u/|\nabla_{\tilde{g}}u|_{\tilde{g}}$ — the unit normal to $\Sigma_{t}$ with respect to $\tilde{g}$;
  • $d\sigma_{\tilde{g}}$ — the area element on $\Sigma_{t}$ induced by $\tilde{g}$;
  • $\alpha_{J}:=\frac{1}{8\pi}K(\eta,\cdot)^{\flat}_{g}$ — the Komar 1-form, using the physical metric $g$ to lower the index.
The angular momentum integral is:
@@
The key difficulty is that the functional $m_{H,J}(t)=\sqrt{m_{H}^{2}(t)+4\pi J(t)^{2}/A(t)}$ involves both $A(t)$ and $J(t)$, and their joint evolution under non-vacuum conditions is not controlled by a simple monotonicity.
@@ 6.4 Monotonicity
$R_{\tilde{g}}\geq 0$ Non-negative integrand Thm 5.8: AM-Lich. ensures this
Vacuum in exterior $J$-conservation ($d^{\dagger}\alpha_{J}=0$) Lem 6.6 Step 6: $\Sigma_{t}\subset M_{\mathrm{ext}}$
Axisymmetry ($\eta$) Komar $J$ well-defined Hypothesis (H2) of Thm 1.2
$A(t)\geq 8\pi|J|$ Sub-extremality factor $\geq 0$ Thm 7.2: preserved by flow
Level set regularity Integration well-defined Rem 6.4: $C^{1,\beta}$ by Tolksdorf
Homologous $\Sigma_{t}$ Stokes for $J$-conservation Lem 6.6: level sets cobordant
@@
where $m_{H}(t)=\sqrt{A(t)/(16\pi)}(1-W(t)/16\pi)$ is the standard Hawking mass, $A(t)=|\Sigma_{t}|_{\tilde{g}}$ is the area, $W(t)=\int_{\Sigma_{t}}H^{2}\,dA_{\tilde{g}}$ is the Willmore functional, and $J$ is the conserved Komar angular momentum.
@@
    Area bound: $A^{\prime}=\displaystyle\int_{\Sigma_{t}}\frac{H}{|\nabla u|}\,d\sigma$ (Step 8c)
@@
    Sub-extremality factor: $1-(8\pi|J|/A)^{2}\geq 0$ when $A\geq 8\pi|J|$ (Step 8g)
@@
    Final bound: $\frac{d}{dt}m_{H,J}^{2}\geq\frac{1}{8\pi}\int_{\Sigma_{t}}\frac{R_{\tilde{g}}+2|\mathring{h}|^{2}}{|\nabla u|}\left(1-\frac{64\pi^{2}J^{2}}{A^{2}}\right)d\sigma\geq 0$ (Step 8h)
@@
Define the Willmore deficit $W(t):=\frac{1}{16\pi}\int_{\Sigma_{t}}H^{2}\,d\sigma$, so $m_{H}=\sqrt{A/(16\pi)}(1-W)$ and $m_{H}^{2}=\frac{A}{16\pi}(1-W)^{2}$.
@@
    Uniform-in-$p$ estimates (Lemma 6.39): Tolksdorf–Lieberman–DiBenedetto regularity with constants independent of $p\in(1,2]$.
@@
$\int_{a}^{b}\mathcal{F}_{p}(t)\,dt\leq\int_{a}^{b}\int_{\Sigma_{t}^{(p)}}\frac{C}{c_{0}}\,dA\,dt\leq\frac{C}{c_{0}}\mathrm{Vol}(\{a\leq u_{p}\leq b\})\cdot\sup_{t}A_{p}(t)^{-1}\cdot(\text{geom.\ bounds}).$
@@
    Monotonicity constant: The monotonicity formula coefficient $\frac{d}{dt}m_{H,J}^{2}(t)\geq 0$ holds with explicit lower bound:
@@
Conclusion: The “a.e. tt” condition is technically necessary for the pointwise derivative formula, but global monotonicity $m_{H,J}(1)\geq m_{H,J}(0)$ holds unconditionally by integration.
@@
    By absolute continuity of $m_{H,J}(t)$ (following from absolute continuity of $m_{H}(t)$ and $A(t)$), the a.e. derivative condition $\frac{d}{dt}m_{H,J}(t)\geq 0$ implies $m_{H,J}(t_{2})\geq m_{H,J}(t_{1})$ for all $t_{1}<t_{2}$.
@@
$\leq C\|\tilde{g}_{\epsilon}-\tilde{g}\|_{C^{0}}\cdot A_{p,\epsilon}(t)+C\|\nabla(u_{p,\epsilon}-u_{p})\|_{C^{0}}\cdot\text{Perimeter}(\Sigma_{t})$
@@
$|\mathcal{M}_{p,J,\epsilon}(t)-\mathcal{M}_{p,J}(t)|\leq C(A_{0},J)|A_{p,\epsilon}(t)-A_{p}(t)|\leq C\epsilon^{\min(\beta_{0},1)}.$
@@
    Explicit Moore–Osgood verification: For the double limit $\lim_{p\to 1^{+}}\lim_{\epsilon\to 0^{+}}A_{p,\epsilon}(t)=\lim_{\epsilon\to 0^{+}}\lim_{p\to 1^{+}}A_{p,\epsilon}(t)$, we verify Moore–Osgood hypotheses explicitly:
@@
    • Pointwise limit existence: $\lim_{p\to 1^{+}}A_{p}(t)$ exists by $W^{1,1}$-compactness of $\{u_{p}\}$.
    • Quantitative uniformity: Setting $\epsilon(p)=(p-1)^{1/\beta_{0}}$ yields $|A_{p,\epsilon(p)}(t)-A_{1}(t)|\leq C(p-1)^{\min(1,\gamma)}$.

### v4

@@ 6.2 The AM-AMO Functional
$m_{H,J}(t):=\sqrt{m_{H}^{2}(t)+\frac{4\pi J^{2}}{A(t)}}=\sqrt{\frac{A(t)}{16\pi}\left(1-W(t)\right)^{2}+\frac{4\pi J^{2}}{A(t)}},$ (30)
@@
For sub-extremal surfaces with $A(t)\geq 8\pi|J|$ (ensured by Theorem 7.2), the argument of the outer square root is non-negative. The Willmore bound $W(t)\leq 1$ is established in Lemma 6.10 below.
@@
$\frac{1}{16\pi}\int_{\Sigma_{t}}H^{2}\,dA\leq 1\quad\Longleftrightarrow\quad W(t)\leq 1.$
Step 8: Verification at boundary and infinity. At $t=0$ (the MOTS), the surface $\Sigma$ has $H_{\tilde{g}}|_{\Sigma}=0$ (minimal in the conformal metric, see Lemma 8.4), so $W(0)=0<1$. As $t\to 1$ (infinity), large coordinate spheres of radius $R$ have $H\approx 2/R$ and area $\approx 4\pi R^{2}$, giving $W\approx\frac{1}{16\pi}(4/R^{2})(4\pi R^{2})=1-O(1/R)$. For intermediate $t$, the monotonicity $m_{H}^{\prime}(t)\geq 0$ implies (1W(t))(1-W(t)) remains non-negative throughout.
The bound $W(t)\leq 1$ is therefore a consequence of the AMO monotonicity framework, not an independent assumption. ∎
@@
We adopt the convention that $W(t)=\frac{1}{16\pi}\int_{\Sigma_{t}}H^{2}\,dA$ is the normalized Willmore energy (dimensionless). This factor equals 0 when $H\equiv 0$ (minimal surfaces), lies in $[0,1]$ for topological 2-spheres with non-negative scalar curvature ambient manifolds, and approaches $1^{-}$ as $t\to 1$ (large coordinate spheres). This convention simplifies the Hawking mass formula to $m_{H}=\sqrt{A/(16\pi)}(1-W)$. We consistently use this normalized definition throughout the paper to avoid confusion with the unnormalized energy $\int H^{2}dA$.
@@
$W(t)=\frac{1}{16\pi}\int_{\Sigma_{t}}H^{2}\,d\sigma\approx\frac{1}{16\pi}\cdot 4\pi R^{2}\cdot\frac{4}{R^{2}}=1-O(R^{-1}).$
@@ 6.3 Angular Momentum Conservation
Answer: The key insight is the separation of roles. The underlying smooth manifold is the same: $M=\bar{M}=\tilde{M}$ as topological spaces (the Jang and conformal constructions are diffeomorphisms, not changes of the underlying manifold). The level sets $\Sigma_{t}=\{u=t\}$ are embedded submanifolds of $M$, defined using $\tilde{g}$ but living in the same $M$ where $(g,K)$ are defined. The Komar 1-form $\alpha_{J}=\frac{1}{8\pi}K(\eta,\cdot)^{\flat}_{g}$ is a well-defined 1-form on $M$, independent of any choice of Riemannian metric. The integral $J(\Sigma_{t})=\int_{\Sigma_{t}}\star_{g}\alpha_{J}$ is computed using the Hodge dual with respect to the physical metric $g$, not the conformal metric $\tilde{g}$. Thus $J(\Sigma_{t})$ is a well-defined quantity: integrate the fixed 2-form $\star_{g}\alpha_{J}$ (determined by $(g,K)$) over the surface $\Sigma_{t}$ (located using $\tilde{g}$).
@@
Answer: We do not claim that $d^{\dagger}_{\tilde{g}}\alpha_{J}=0$. Instead, co-closedness is established for the physical metric: $d^{\dagger}_{g}\alpha_{J}=0$. This is equivalent to $d(\star_{g}\alpha_{J})=0$, meaning $\star_{g}\alpha_{J}$ is a closed 2-form. Since the exterior derivative $d$ is metric-independent (a purely topological operation), $d(\star_{g}\alpha_{J})=0$ holds on the smooth manifold $M$ regardless of which metric is used to parametrize surfaces. The conservation law is a consequence of Stokes’ theorem applied to the closed 2-form $\star_{g}\alpha_{J}$, not a conformal invariance statement.
@@
Answer: Yes. The constructions preserve axisymmetry. The Jang equation with axisymmetric boundary conditions yields an axisymmetric solution $f$, so $\mathcal{L}_{\eta}\bar{g}=0$ where $\bar{g}=g+df\otimes df$. The AM-Lichnerowicz equation with axisymmetric data yields an axisymmetric conformal factor $\phi$, so $\mathcal{L}_{\eta}\tilde{g}=0$ where $\tilde{g}=\phi^{4}\bar{g}$. Therefore $\eta$ remains a Killing field for $\tilde{g}$, and the $p$-harmonic flow respects the symmetry. However, this is not needed for conservation: even if $\eta$ were not Killing for $\tilde{g}$, the closed form $\star_{g}\alpha_{J}$ would still have constant flux through homologous surfaces.
@@
Answer: The Jang manifold $\bar{M}$ has a cylindrical end $\mathcal{C}\cong[0,\infty)\times\Sigma$. The Komar 1-form $\alpha_{J}$ extends smoothly to the cylindrical end (it is defined from $(g,K)$, which are smooth), and the 2-form $\star_{g}\alpha_{J}$ is closed throughout $M$, including the cylindrical region. The level sets $\Sigma_{t}$ for $t$ near 0 may approach the MOTS $\Sigma=\Sigma_{0}$, but remain in a region where $\star_{g}\alpha_{J}$ is defined. The flux $\int_{\Sigma_{t}}\star_{g}\alpha_{J}$ is continuous in $t$, even as $t\to 0$, by dominated convergence. The boundary term at the cylindrical end vanishes by the asymptotic analysis in Lemma 5.11.
@@
Theorem 6.14 (Angular Momentum Conservation).
@@
Step (ii): Coordinate calculation. Let $(y^{1},y^{2})$ be local coordinates on $\Sigma$ and extend to coordinates $(y^{1},y^{2},n)$ on $U$ where $n$ is a coordinate transverse to $\Sigma$ with $\Sigma=\{n=0\}$. In these coordinates, the $\gamma$-unit normal is $\nu_{\gamma}=\frac{1}{|\partial_{n}|_{\gamma}}\partial_{n}+(\text{tangential corrections})$, the area element is $d\sigma_{\gamma}=|\partial_{n}|_{\gamma}\sqrt{\det\gamma_{AB}}\,dy^{1}\wedge dy^{2}$ where $\gamma_{AB}$ is the induced metric on $\Sigma$, and the contraction $\alpha_{J}(\nu_{\gamma})=\frac{1}{|\partial_{n}|_{\gamma}}(\alpha_{J})_{n}+(\text{tangential terms})$. The product gives:
@@
The key observation is that the Komar 1-form $\alpha_{J}=\frac{1}{8\pi}K(\eta,\cdot)^{\flat}$ is defined on the physical manifold, but we integrate it over surfaces $\Sigma_{t}$ that are level sets in the conformal picture. This is valid because the underlying smooth manifold $M$ is the same (only the metric changes), the level sets $\Sigma_{t}\subset M$ are well-defined submanifolds independent of which metric we use, the 1-form $\alpha_{J}$ and its exterior derivative $d\alpha_{J}$ are tensorial operations that commute with pullback to any submanifold, and the integral $\int_{\Sigma_{t}}\star_{g}\alpha_{J}$ is computed using the physical metric $g$ for the Hodge dual, making it independent of $\tilde{g}$.
@@
Metric-independence of the Komar integral. We now make explicit which quantities use which metric: $\nu_{\tilde{g}}:=\nabla_{\tilde{g}}u/|\nabla_{\tilde{g}}u|_{\tilde{g}}$ is the unit normal to $\Sigma_{t}$ with respect to $\tilde{g}$; $d\sigma_{\tilde{g}}$ is the area element on $\Sigma_{t}$ induced by $\tilde{g}$; and $\alpha_{J}:=\frac{1}{8\pi}K(\eta,\cdot)^{\flat}_{g}$ is the Komar 1-form, using the physical metric $g$ to lower the index. The angular momentum integral is:
@@
The key difficulty is that the functional
$m_{H,J}(t)=\sqrt{m_{H}^{2}(t)+\frac{4\pi J(t)^{2}}{A(t)}}$
involves both $A(t)$ and $J(t)$, and their joint evolution under non-vacuum conditions is not controlled by a simple monotonicity.
@@ 6.4 Monotonicity
$R_{\tilde{g}}\geq 0$ Non-neg. integrand Thm 5.10: AM-Lich.
Vacuum (ext.) $J$-conservation Lem 6.6: $\Sigma_{t}\subset M_{\mathrm{ext}}$
Axisymmetry Komar $J$ defined (H2) of Theorem 1.1
$A(t)\geq 8\pi|J|$ Sub-extrem. $\geq 0$ Thm 7.2: preserved
Regularity Integrals defined Rem 6.4: $C^{1,\beta}$
Homology Stokes for $J$ Lem 6.6: cobordant
@@
where $m_{H}(t)=\sqrt{A(t)/(16\pi)}(1-W(t))$ is the standard Hawking mass, $A(t)=|\Sigma_{t}|_{\tilde{g}}$ is the area, $W(t)=\frac{1}{16\pi}\int_{\Sigma_{t}}H^{2}\,dA_{\tilde{g}}$ is the Willmore functional, and $J$ is the conserved Komar angular momentum.
@@
    Area bound: $A^{\prime}=\int_{\Sigma_{t}}\frac{H}{|\nabla u|}\,d\sigma$ (Step 8c)
@@
    Sub-extremality: $1-(8\pi|J|/A)^{2}\geq 0$ when $A\geq 8\pi|J|$ (Step 8g)
@@
    Final bound: $\frac{d}{dt}m_{H,J}^{2}\geq\frac{1}{8\pi}\int_{\Sigma_{t}}\frac{R_{\tilde{g}}+2|\mathring{h}|^{2}}{|\nabla u|}\bigl(1-\frac{64\pi^{2}J^{2}}{A^{2}}\bigr)d\sigma\geq 0$ (Step 8h)
@@
Define the Willmore deficit by:
$W(t):=\frac{1}{16\pi}\int_{\Sigma_{t}}H^{2}\,d\sigma.$
Then $m_{H}=\sqrt{A/(16\pi)}(1-W)$ and $m_{H}^{2}=\frac{A}{16\pi}(1-W)^{2}$.
@@
    Uniform-in-$p$ estimates (Lemma 6.39): Tolksdorf–Lieberman regularity with constants independent of $p\in(1,2]$.
@@
$\int_{a}^{b}\mathcal{F}_{p}(t)\,dt$ $\leq\int_{a}^{b}\int_{\Sigma_{t}^{(p)}}\frac{C}{c_{0}}\,dA\,dt$
$\leq\frac{C}{c_{0}}\mathrm{Vol}(\{a\leq u_{p}\leq b\})\cdot\sup_{t}A_{p}(t)^{-1}\cdot(\text{geom.\ bounds}).$
@@
    Monotonicity coefficient: The monotonicity formula coefficient $\frac{d}{dt}m_{H,J}^{2}(t)\geq 0$ holds with explicit lower bound:
@@
Conclusion: The “a.e. tt” condition is technically necessary for the pointwise derivative formula, but global monotonicity $m_{H,J}(1)\geq m_{H,J}(0)$ holds unconditionally via integration of the a.e. non-negative derivative.
@@
The following observations justify our treatment of critical points:
@@
    By absolute continuity, the a.e. derivative condition $\frac{d}{dt}m_{H,J}(t)\geq 0$ implies $m_{H,J}(t_{2})\geq m_{H,J}(t_{1})$ for all $t_{1}<t_{2}$.
@@
$\leq C\|\tilde{g}_{\epsilon}-\tilde{g}\|_{C^{0}}\cdot A_{p,\epsilon}(t)$
$\quad+C\|\nabla(u_{p,\epsilon}-u_{p})\|_{C^{0}}\cdot\text{Perimeter}(\Sigma_{t})$
@@
$|\mathcal{M}_{p,J,\epsilon}(t)-\mathcal{M}_{p,J}(t)|$ $\leq C(A_{0},J)|A_{p,\epsilon}(t)-A_{p}(t)|$
$\leq C\epsilon^{\min(\beta_{0},1)}.$
@@
    Moore–Osgood verification: For the double limit
    $\lim_{p\to 1^{+}}\lim_{\epsilon\to 0^{+}}A_{p,\epsilon}(t)=\lim_{\epsilon\to 0^{+}}\lim_{p\to 1^{+}}A_{p,\epsilon}(t),$
    we verify Moore–Osgood hypotheses explicitly:
@@
    • Pointwise limit: $\lim_{p\to 1^{+}}A_{p}(t)$ exists by $W^{1,1}$-compactness of $\{u_{p}\}$.
    • Uniformity: Setting $\epsilon(p)=(p-1)^{1/\beta_{0}}$ yields $|A_{p,\epsilon(p)}(t)-A_{1}(t)|\leq C(p-1)^{\min(1,\gamma)}$.